To three decimal places, find the value of the first positive x-intercept for the function f(x) = –2cos(x + 1).
ANSWER CHOICES:
0.571
1.571<--- MY ANSWER
–1.081
–2.571

Answer by srinivas.g(540)

f(x) should be to to find x-intercept
it is possible only if x+1 =pi/2
x+1 =1.5707
x=0.5707
result : 0.5707 (0.571)